OpenStudy (anonymous):

What three lengths could be lengths of the sides of a triangle? A) 9cm, 22cm, 11cm B) 21cm, 7cm, 6cm C) 12cm, 5cm, 17cm or D) 10cm,15cm, 24cm

OpenStudy (tkhunny):

Pick two sides. Add them. Is the sum greater than the length of the third side? Example: 9, 22, 11 9+22 = 31 > 11 Good 11+22 = 33 > 9 Good 9+11 = 20 < 22 No Good! Okay, move on from that one. That cannot be a triangle.

OpenStudy (anonymous):

oh okay I get it now! Thank you so much!!

OpenStudy (tkhunny):

It's calle the "Triangle Inequality". It is your friend!